OpenStudy (hydrogen):

Fan and medal! Use the Law of Conservation of Mass to calculate the expected mass of the product in the following reaction if 120 g of calcium reacts with 96 g of oxygen gas:

OpenStudy (hydrogen):

Ca + O2 --> CaO I am to write a sentence to explain my answer. I assumed the reaction had to be balanced, so I came up with: 2Ca + O2 -> 2CaO However, I can't seem to get past this. The Law of Conservation of Mass states that mass does not change during a chemical reaction, and the mass of the products is equal to the mass of the reactants. But, what do I write for this answer? I'm unsure of how to figure out the expected mass. This is 10th grade Science. Thank you very much for any help!! I'd really appreciate it!

OpenStudy (alphadxg):

First thing to do is to find which among Ca and O2 is the limiting reactant. Solve for theoretical mole ratio based on the given chemical balanced equation: 2Ca + O2 =====> 2CaO TMR: Ca / O2 = 2/1 = 2 Solve for actual mole ratio by dividing moles of Ca and O2 from the given masses: AMR: Ca / O2 = (120 g / 40 g/mol) / (96 g / 32 g/mol) = 1 Compare TMR and AMR: If AMR < TMR it means that your limiting reactant is in the numerator. Otherwise if AMR > TMR, then your limiting reactant is in the denominator. In this particular problem, AMR < TMR, so the limiting reactant is Ca. Using moles of Ca, we can calculate mass of CaO. mass of CaO = (mass of Ca / MW of Ca) x (1 mole CaO / 1 mole Ca) x MW of CaO mass of CaO = (120 g / 40 g/mol) x ( 1/1 ) x 56 g/mol mass of CaO = 168 grams ====> answer

OpenStudy (hydrogen):

Thank you very much for your help!!